OpenStudy (ilovebmth1234):

When the [H+] in a solution is 1.7 × 10^-9 M, what is the pOH ?

OpenStudy (ilovebmth1234):

A) 8.80 B) 5.20 C) 14 D) 8.60

OpenStudy (ilovebmth1234):

@james1107

OpenStudy (james1107):

@Lexi.babyy

OpenStudy (james1107):

Aug 4, 2011 - Best Answer: 14 = pH + pOH -log(1.7x10^-9) = pH =(8.77) Calculate the [OH ‾ ], [H+], and the pH of 0.17 M solutions of each of the following ...

OpenStudy (james1107):

according to this it would be a

OpenStudy (ilovebmth1234):

so C according to that?

OpenStudy (ilovebmth1234):

No 14 isnt the answer but thats okay :3 thank you

OpenStudy (james1107):

look at everything in the parentheses (8.77)

OpenStudy (rushwr):

Do it as i said earlier ? @ilovebmth1234

OpenStudy (ilovebmth1234):

ohhh lol and yeah but its kinda confusing like i said

OpenStudy (james1107):

did you see what i ws talking about

OpenStudy (james1107):

Aug 4, 2011 - Best Answer: 14 = pH + pOH -log(1.7x10^-9) = pH = 8.77 ... Calculate the [OH ‾ ], [H+], and the pH of 0.17 M solutions of each of the following ...

OpenStudy (james1107):

pH=8.77

OpenStudy (ilovebmth1234):

when you add them together how would i do that exactly?

OpenStudy (james1107):

type in just the question so you can looh at the first site that comes up i think its like answers .com or something its the very first website

OpenStudy (james1107):

look at the

OpenStudy (ilovebmth1234):

can you take a screen shot and post it? i cant get on yahoo its blocked :/

OpenStudy (james1107):

how do you do that on a laptop

OpenStudy (ilovebmth1234):

press your start button click all programs and go to accessories and find snip tool and click on it

OpenStudy (ilovebmth1234):

or try control+ alt + prnt scrn(aka- end)

OpenStudy (rushwr):

Do u want me to go thru it again ? @ilovebmth1234

OpenStudy (ilovebmth1234):

if you make it to where i can understand it

OpenStudy (rushwr):

okai

OpenStudy (rushwr):

When the concentration of H^+ is given we can't directly use an equation to find the pOH To find we need to know the pH of the solution.

OpenStudy (rushwr):

\[pH = - \log[H ^{+}]\] \[pH= -\log[1.7 * 10^{-9}]\]

OpenStudy (rushwr):

find that value first !

OpenStudy (rushwr):

@ilovebmth1234 hey are u working it out ?

OpenStudy (ilovebmth1234):

okay im trying to

OpenStudy (rushwr):

I'll make it easy for u -log10^-9 = 9 -log1.7 u need to find the log value for 1.7 and put a minus charge in front !

OpenStudy (ilovebmth1234):

what does -log mean?

OpenStudy (rushwr):

log values !!!!!!!!!!!!!!!!!!!!!!!! -log is the minus log value

OpenStudy (ilovebmth1234):

damn my bad sorry i dont quit understand this like you do v.v

OpenStudy (rushwr):

lol so did u find the value now ?

OpenStudy (rushwr):

\[-\log10^{-9} = 9\] \[-\log1.7 = - 0.23\] \[pH = 9 - 0.23 = 8.77\]

OpenStudy (ilovebmth1234):

A

OpenStudy (ilovebmth1234):

or D

OpenStudy (rushwr):

pH + pOH = 14 Since pH = 8.77 pOH = 14 - pH = 14 - 8.77 pOH = 5.23

OpenStudy (rushwr):

pH is not pOH

OpenStudy (ilovebmth1234):

right right

OpenStudy (ilovebmth1234):

so B would be the answer

OpenStudy (rushwr):

yes yes

OpenStudy (ilovebmth1234):

great thanks :3 could you give @james1107 a medal for his help too?

OpenStudy (rushwr):

ah yes sure :)